I need to determine if the statement a < b, then a + c < b + c is true or false. Thank you, Bonnie
Found 2 solutions by rothauserc, ramkikk66:
Answer by rothauserc(4718)

a < b, then a + c < b + c is true or false.....this inequality is true if and only if c > or = 0

Answer by ramkikk66(644)

If a < b, a + c < b + c
This inequality is true for all real values of a,b,c.
Even if c is negative, if a < b, a + c will be less than b + c.
Example: Check with a = 7, b = 10, c = -20
7 < 10
a + c = 7 - 20 = -13
b + c = 10 - 20 = -10
-13 < -10
:)